Could use some help here, did I do this correctly. Thanks
If f(x)=3x-1 and g(x)=2X-1
Find f(g(2))
3x(2x-1(2))-1
3x(4x-2)-1
12x^-6x-1

Answer by Earlsdon(6103)

If f(x) = 3x-1 and g(x) = 2x-1, find f(g(2))
Start with g(2) = 2(2)-1 = 4-1 = 3
Now do f(3) = 3(3)-1 = 9-1 = 8